# Angular integrals with three denominators via IBP, mass reduction, dimensional shift, and differential equations

### v1

@@ Abstract
I Introduction
Angular integrals [1, 2, 3, 4, 5, 6, 7] are an elementary building block of phase-space integrals arising in perturbative quantum field theory [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In the Quantum Chromodynamics (QCD) literature they have been used for example in theoretical predictions for the Drell-Yan process (DY) [24, 25, 26, 27, 14], deep-inelastic scattering (DIS) [28, 29], semi-inclusive deep-inelastic scattering (SIDIS) [30, 31], prompt-photon production [32, 33], hadron-hadron scattering [34], heavy quark production [3], and single-spin asymmetries [35, 36]. In processes with massless particles, collinear singularities are present, necessitating a calculation in $d=4-2\varepsilon$ dimensions [37, 38].
@@ I Introduction
While the two-denominator case is well understood with analytic results in $d$ dimensions [4, 5] as well as $\varepsilon$-expansions to all orders [5], far less is known in the case of more denominators. In [4], the massless three-denominator integral has been calculated to order $\varepsilon^{0}$ and a general Mellin-Barnes representation for angular integrals was given. Recently, multi-denominator angular integrals have been studied in the limit of small masses in [7]. The $\varepsilon$-expansions for the three-denominator integral with an arbitrary number of masses were given to order $\varepsilon^{0}$ to leading power in the masses. Further showcasing the newly increased interest in the subject, the $\varepsilon$-expansion of the single-massive four-denominator angular integral has been studied at as an example for a novel regularization scheme with connections to tropical geometry [39].
@@
Following the notation of [4, 5, 7], the integral family under consideration is
$I_{j_{1}j_{2}j_{3}}^{(m)}(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33};\varepsilon)=\int\frac{\mathrm{d}\Omega_{d-1}(k)}{\Omega_{d-3}}\frac{1}{(v_{1}\cdot k)^{j_{1}}(v_{2}\cdot k)^{j_{2}}(v_{3}\cdot k)^{j_{3}}}\,,$ (1)
@@
$k$ $=(1,\dots,\sin\theta_{1}\sin\theta_{2}\cos\theta_{3},\sin\theta_{1}\cos\theta_{2},\cos\theta_{1})\,,$
$v_{1}$ $=(1,\mathbf{0}_{d-2},\beta_{1})\,,$
$v_{2}$ $=(1,\mathbf{0}_{d-3},\beta_{2}\sin\chi_{2}^{(1)},\beta_{2}\cos\chi_{2}^{(1)})\,,$
$v_{3}$ $=(1,\mathbf{0}_{d-4},\beta_{3}\sin\chi_{3}^{(1)}\sin\chi_{3}^{(2)},\beta_{3}\sin\chi_{3}^{(1)}\cos\chi_{3}^{(2)},\beta_{3}\cos\chi_{3}^{(1)})\,,$
@@
where $\Omega_{d}=2\pi^{d/2}/\Gamma(d/2)$. The normalization is chosen in accordance with [4] to remove factors of the Euler-Mascheroni constant $\gamma_{\text{E}}$ from the $\varepsilon$-expansion. The integral depends on the invariants $v_{ij}\equiv v_{i}\cdot v_{j}$. The denominator powers $j_{1}$, $j_{2}$, $j_{3}$ are assumed to be integers in the following. The superscript $m=0,1,2,3$ characterizes the number of non-zero masses $v_{11}$, $v_{22},v_{33}$.
For the calculation of the $\varepsilon$-expansion we use the method of differential equations which is well known for the calculation of loop integrals [40, 41, 42, 43, 44, 45]. Generally, the strategy consists of the following steps:
@@
    Derive recursion relations from integration-by-parts (IBP) identities to express every $I_{j_{1}j_{2}j_{3}}$ in a basis of master integrals.
@@
In the particular case at hand, the structure of the angular integrals allows for several simplifications to this general approach. Most importantly, by studying the behavior under dimensional shift $d\rightarrow d+2$ it is possible to reconstruct the pole and finite part of the $\varepsilon$-expansion solely in terms of known two-denominator integrals. Simultaneously, the dimensional shift can be used for analytic continuation to obtain finite integral representations – fit for numerical checks – for each order in the expansion. To facilitate keeping an overview of what follows, figure 1 provides a graphical overview of the steps taken to extract the $\varepsilon$-expansion of the general three-denominator angular integral with $m$ masses $I_{j_{1}j_{2}j_{3}}^{(m)}$.
@@
Figure 1: This flowchart provides an overview of the calculation of the $\varepsilon$-expansion of the general three-denominator angular integral $I_{j_{1}j_{2}j_{3}}^{(m)}$. In a first step, recursion relations (see app. C) derived from IBP relations (see sec. II) are used for a reduction to the master integral $I_{111}^{(m)}$. In a second step, the double- and triple-massive integrals are expressed in terms of massless and single-massive ones through mass reduction formulae derived from the two-point splitting lemma (see sec. III). In a third step, a combination of a dimensional-shift identity, relating integrals in $d$ and $d+2$ dimensions, with the recursion relations allows for the determination of pole and finite part in terms of known two-denominator integrals (see sec. IV). In a final step, the order $\varepsilon$ contribution is calculated by applying the method of differential equations – requiring suitable differential operators for angular integrals (see sec. V) and again making use of the recursion relations – to the massless and single-massive master integral in $d=6-2\varepsilon$ dimensions (see sec. VI and VII). Results are found in sec. VIII. Image created with JaxoDraw [46].
@@
The remainder of the paper is organized as follows. In section II, we derive IBP relations to reduce $I_{j_{1},j_{2},j_{3}}$ to master integrals. Section III recalls the two-point splitting lemma, which allows us to express double- and triple-massive integrals in terms of the single-massive integral. The dimensional-shift identity introduced in section IV allows us to express the pole and finite part of the three-denominator integral solely in terms of known two-denominator integrals. In section V, we determine differential operators, which we subsequently use to derive differential equations for the master integrals in section VI. The integration of these differential equations is discussed in section VII. Results for the massless, single-massive, double-massive, and triple-massive master integrals $I^{(0,1,2,3)}_{111}$ are given in section VIII. The generalization towards angular integrals with more than three denominators is briefly discussed in section IX, before we conclude in section X. Appendix A lists the scalar products used for the mass reduction formula, while appendices B and C give the full form of the IBP relations and the recursion relations obtained from the latter. In appendix D, a proof of the general dimensional-shift identity for angular integrals is provided. Appendix E lists the expansions for the dimensionally-shifted two-denominator integrals that are used as input for the calculation of the three-denominator master integrals. The explicit calculation of the boundary value for the massless master integral is given in appendix F. Integrals useful for the calculation of the master integrals are found in appendix G, and appendix H recalls the Clausen function appearing in the order $\varepsilon$ of the three-denominator angular integrals. The ancillary Mathematica file provides the function AngularIntegral[{j1_,j2_,j3_},{v12_,v13_,v23_,v11_,v22_,v33_}] which evaluates the three-denominator angular integral in $d=4-2\varepsilon$ dimensions up to order $\varepsilon$.
II IBP relations
@@
Since they are the foundation for everything that follows, we continue by deriving IBP relations to reduce $I_{j_{1},j_{2},j_{3}}$ to master integrals. IBP relations for angular integrals are derived in a similar fashion to loop integrals. The only difference is that permissible vectors in the IBP relation have to be tangential to the spherical integration surface. Concretely, the IBPs read
@@
where the vector $\vec{\xi}$ must satisfy $\vec{\xi}\cdot\vec{k}\overset{!}{=}0$.
@@
For general $j_{1,2,3}$ these IBP relations can be combined into identities that only either raise or lower the sum of indices. This allows to bring them in the form of recursion relations that systematically reduce all integrals $I_{j_{1},j_{2},j_{3}}$ into the master integrals $I_{111}$, $I_{110}$, $I_{100}$, $I_{000}$ (+ permutation of indices), bypassing the need for Laporta’s algorithm [47]. The explicit recursion relations are given in appendix C.
To calculate the master integrals, we use the method of differential equations. For this, we write the master integrals in the form of eq. (1) and differentiate both sides with respect to the kinematic invariants $v_{ij}$, which will lead to new angular integrals on the right-hand side – details of the differentiation are discussed in section V. By reducing the right-hand side to master integrals via the recursion relations, we obtain a system of differential equations for the master integrals. For a pedagogical introduction to the differential equation approach in the context of loop integrals see [44, 45].
Since the $\varepsilon$-expansion of two-denominator angular integrals is known [3, 4, 5], the only unknown master integral in our case is $I_{111}$. By explicitly inserting the known integrals into the coupled system of differential equations, it is reduced to one closed differential equation for $I_{111}$.
@@
    By two-mass splitting we can always reduce the triple-massive integral to single-massive integrals. We will be discussing this next in section III. This simplifies the differential equation by reducing the number of variables. The number of master integrals is further reduced since the massless one-denominator integral $I_{100}$ is reducible to $I_{000}$.
@@
    By using a dimensional-shift identity $d\rightarrow d+2$ we obtain the pole and – somewhat surprisingly – also the finite part of the three-denominator integral “for free”, i.e. in terms of known integrals with fewer denominators. We will look at this in section IV.
@@
    By symmetry it suffices to look at a single derivative, say ${\partial}/{\partial v_{12}}$, for the massless case. From there we only need a single boundary value for which we choose the symmetric point $v_{12}=v_{13}=v_{23}=1$. To construct the single-massive integral from the differential equation with respect to ${\partial}/{\partial v_{11}}$ we can use the previously calculated massless integral as a boundary condition. This will be done in sections VI and VII.
III Mass reduction via two-point splitting
It is a truth universally acknowledged, that a single-massive angular integral is sufficient to calculate multi-mass angular integrals. This is possible due to the two-point splitting lemma [5, 6], which states that for any two vectors $v_{1}$ and $v_{2}$, we can choose any scalar $\lambda_{(12)}$ and construct the linear combination ${v_{(12)}\equiv(1-\lambda_{(12)})\,v_{1}+\lambda_{(12)}\,v_{2}}$ to obtain
@@ III Mass reduction via two-point splitting
If we consider two massive vectors $v_{1}$ and $v_{2}$ and choose $\lambda_{(12)}$ such that $v_{(12)}$ is massless, i.e. $v_{(12)(12)}=0$, eq.(9) splits the product of two massive denominators into the sum of two single-massive denominator products. Iterative use of this identity on pairs of massive denominators allows for a reduction to the single-massive case. Hence angular integrals with an arbitrary number of masses can always be expressed through at most single-massive integrals.  [^1] A quite similar idea of splitting loop integrals into integrals with fewer masses was already put forward a long time ago by ’t Hooft and Veltman [48].
@@
with $X_{12}\equiv v_{12}^{2}-v_{11}v_{22}$. Note that there is the geometrical interpretation $X_{12}=4\,\mathbb{V}(\vec{v}_{1},\vec{v}_{2})^{2}+(\vec{v}_{1}-\vec{v}_{2})^{2}$, where $\mathbb{V}(\vec{v}_{1},\vec{v}_{2})$ denotes the area of the triangle spanned by $\vec{v}_{1}$ and $\vec{v}_{2}$. Hence, for real vectors $\vec{v}_{1,2}$, $X_{12}$ is non-negative and vanishes if and only if $\vec{v}_{1}=\vec{v}_{2}$. The scalar products of the auxiliary vectors are ${v_{(12)(12)}=v_{(21)(21)}=0}$, ${v_{1(12)}=-\lambda_{(12)}\sqrt{X_{12}}}$, ${v_{2(21)}=-\lambda_{(21)}\sqrt{X_{12}}}$, and ${v_{(12)(21)}=2X_{12}/(2v_{12}-v_{11}-v_{22})}$. Note that $v_{1(12)}$ respectively $v_{2(21)}$ vanish in the massless limits $v_{11}\rightarrow 0$ respectively $v_{22}\rightarrow 0$ Using these scalar products, a compact form of the two-mass splitting is
@@
$I_{111}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})=\frac{1}{\sqrt{X_{12}}}$ $\left[v_{(12)(21)}I_{111}^{(0)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)}\right)-v_{1(12)}I_{111}^{(1)}\left(v_{13},v_{1(12)},v_{3(12)},v_{11}\right)\right..$
$\left.-v_{2(21)}I_{111}^{(1)}\left(v_{23},v_{2(21)},v_{3(21)},v_{22}\right)\right]$ (14)
@@
$I_{111}^{(3)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33})=\frac{1}{\sqrt{X_{12}}}\Big{\{}v_{(12)(21)}I_{111}^{(1)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)},v_{33}\right)$
$-\frac{v_{1(12)}}{\sqrt{X_{13}}}\Big{[}v_{(13)(31)}I_{111}^{(0)}\left(v_{(12)(13)},v_{(12)(31)},v_{(13)(31)}\right)-v_{1(13)}I_{111}^{(1)}\left(v_{1(12)},v_{1(13)},v_{(12)(13)},v_{11}\right)$
$\qquad-v_{3(31)}I_{111}^{(1)}\left(v_{3(12)},v_{3(31)},v_{(12)(31)},v_{33}\right)\Big{]}$
$-\frac{v_{2(21)}}{\sqrt{X_{23}}}\Big{[}v_{(23)(32)}I_{111}^{(0)}\left(v_{(21)(23)},v_{(21)(32)},v_{(23)(32)}\right)-v_{2(23)}I_{111}^{(1)}\left(v_{2(21)},v_{2(23)},v_{(21)(23)},v_{22}\right)$
$\qquad-v_{3(32)}I_{111}^{(1)}\left(v_{3(21)},v_{3(32)},v_{(21)(32)},v_{33}\right)\Big{]}\Big{\}}\,,$ (15)
@@
$\tilde{I}^{(m)}_{111}(v_{ij},v_{ik},v_{jk},v_{ii},v_{jj},v_{kk})\equiv\sqrt{Y_{ijk}}\,I^{(m)}_{111}(v_{ij},v_{ik},v_{jk},v_{ii},v_{jj},v_{kk})\,,$ (20)
@@
$\tilde{I}_{111}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})=\tilde{I}_{111}^{(0)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)}\right)-\tilde{I}_{111}^{(1)}\left(v_{13},v_{1(12)},v_{3(12)},v_{11}\right)-\tilde{I}_{111}^{(1)}\left(v_{23},v_{2(21)},v_{3(21)},v_{22}\right),$ (21)

### v4

# Auf der Morgenstelle 14, 72076 Tübingen, Germany

@@ Abstract
preprint: arXiv:2410.18177
1 Introduction
Angular integrals Schellekens:1981 ; vanNeerven:1985 ; Beenakker:1988 ; Somogyi:2011 ; Lyubovitskij:2021 ; Wunder:2024 ; Smirnov:2024 are an elementary building block of phase-space integrals arising in perturbative quantum field theory Bolzoni:2010 ; Anastasiou:2013 ; Lillard:2016 ; Kotlarski:2016 ; Lionetti:2018 ; Specchia:2018 ; Bahjat-Abbas:2018 ; Baranowski:2020 ; Blumlein:2020 ; Isidori:2020 ; Alioli:2022 ; Assi:2023 ; Catani:2023 ; Pal:2023 ; Devoto:2024 ; Rowe:2024 . In the Quantum Chromodynamics (QCD) literature they have been used for example in theoretical predictions for the Drell-Yan process (DY) Matsuura:1989 ; Matsuura:1990 ; Hamberg:1991 ; Mirkes:1992 ; Bahjat-Abbas:2018 , deep-inelastic scattering (DIS) Duke:1982 ; Hekhorn:2019 , semi-inclusive deep-inelastic scattering (SIDIS) Anderle:2016 ; Wang:2019 , prompt-photon production Gordon:1993 ; Coriano:1996 ; Rein:2024 , hadron-hadron scattering Ellis:1980 , heavy quark production Beenakker:1988 , and single-spin asymmetries Schlegel:2012 ; Ringer:2015 . In processes with massless particles, collinear singularities are present, necessitating a calculation in $d=4-2\varepsilon$ dimensions tHooft:1972 ; Bollini:1972 .
@@ 1 Introduction
While the two-denominator case is well understood with analytic results in $d$ dimensions Somogyi:2011 ; Lyubovitskij:2021 as well as $\varepsilon$-expansions to all orders Lyubovitskij:2021 , far less is known in the case of more denominators. In Somogyi:2011 , the massless three-denominator integral has been calculated to order $\varepsilon^{0}$ and a general Mellin-Barnes representation for angular integrals was given. Recently, multi-denominator angular integrals have been studied in the limit of small masses in Smirnov:2024 . The $\varepsilon$-expansions for the three-denominator integral with an arbitrary number of masses were given to order $\varepsilon^{0}$ to leading power in the masses. Further showcasing the newly increased interest in the subject, the $\varepsilon$-expansion of the single-massive four-denominator angular integral has been studied at as an example for a novel regularization scheme with connections to tropical geometry Salvatori:2024 .
@@
Following the notation of Somogyi:2011 ; Lyubovitskij:2021 ; Smirnov:2024 , the integral family under consideration is
$I_{j_{1},j_{2},j_{3}}^{(m)}(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33};\varepsilon)=\int\frac{\mathrm{d}\Omega_{d-1}(k)}{\Omega_{d-3}}\frac{1}{(v_{1}\cdot k)^{j_{1}}(v_{2}\cdot k)^{j_{2}}(v_{3}\cdot k)^{j_{3}}}\,$, (1)
@@
$k$ $={\color[rgb]{0,0,0}(1,\vec{k})=}(1,\dots,\sin\theta_{1}\sin\theta_{2}\cos\theta_{3},\sin\theta_{1}\cos\theta_{2},\cos\theta_{1})\,$,
$v_{1}$ $={\color[rgb]{0,0,0}(1,\vec{v}_{1})=}(1,\vec{0}_{d-2},\beta_{1})\,$,
$v_{2}$ $={\color[rgb]{0,0,0}(1,\vec{v}_{2})=}(1,\vec{0}_{d-3},\beta_{2}\sin\chi_{2}^{(1)},\beta_{2}\cos\chi_{2}^{(1)})\,$,
$v_{3}$ $={\color[rgb]{0,0,0}(1,\vec{v}_{3})=}(1,\vec{0}_{d-4},\beta_{3}\sin\chi_{3}^{(1)}\sin\chi_{3}^{(2)},\beta_{3}\sin\chi_{3}^{(1)}\cos\chi_{3}^{(2)},\beta_{3}\cos\chi_{3}^{(1)})\,$,
@@
where $\Omega_{d}=2\pi^{d/2}/\Gamma(d/2)$. The normalization is chosen in accordance with Somogyi:2011 to remove factors of the Euler-Mascheroni constant $\gamma_{\text{E}}$ from the $\varepsilon$-expansion. Here and in the following the spatial part of a $d$-vector $\xi$ is indicated by an arrow as $\vec{\xi}$. The dot product between two $d$-vectors denotes the usual Minkowski scalar product, between spatial vectors it means the Euclidean scalar product, i.e. for normalized $d$-vectors $\xi$ and $\eta$ it is $\xi\cdot\eta=1-\vec{\xi}\cdot\vec{\eta}$. The integral depends on the invariants $v_{ij}\equiv v_{i}\cdot v_{j}$. The parametric variables $\beta_{i}$ and $\chi_{j}^{(i)}$, where $1\leq i\leq 3$ and $i<j\leq 3$, can be expressed through the invariant variables $v_{ij}$ via
$\beta_{i}$ $=\sqrt{1-v_{ii}}\,,\;\cos\chi_{j}^{(1)}=\frac{1-v_{1j}}{\beta_{1}\beta_{j}}\,,\;\cos\chi_{j}^{(2)}=\frac{1-v_{2j}-\beta_{2}\beta_{j}\cos\chi_{2}^{(1)}\cos\chi_{j}^{(1)}}{\beta_{2}\beta_{j}\sin\chi_{2}^{(1)}\sin\chi_{j}^{(1)}}\,$,
$\cos\chi_{j}^{(3)}$ $=\frac{1-v_{3j}-\beta_{3}\beta_{j}\left(\cos\chi_{3}^{(1)}\cos\chi_{j}^{(1)}+\sin\chi_{3}^{(1)}\sin\chi_{j}^{(1)}\cos\chi_{3}^{(2)}\cos\chi_{j}^{(2)}\right)}{\beta_{3}\beta_{j}\sin\chi_{3}^{(1)}\sin\chi_{j}^{(1)}\sin\chi_{3}^{(2)}\sin\chi_{j}^{(2)}}\,.$ (3)
The denominator powers $j_{1}$, $j_{2}$, $j_{3}$ are assumed to be integers in the following. The superscript $m=0,1,2,3$ characterizes the number of non-zero masses $v_{11}$, $v_{22},v_{33}$.
For the calculation of the $\varepsilon$-expansion we use the method of differential equations which is well known for the calculation of loop integrals Kotikov:1990 ; Remiddi:1997 ; Gehrmann:1999 ; Henn:2013 ; Henn:2014 ; Badger:2023 . Generally, the strategy consists of the following steps:
@@
    Derive recursion relations from integration-by-parts (IBP) identities to express every $I_{j_{1},j_{2},j_{3}}$ in a basis of master integrals.
@@
In the particular case at hand, the structure of the angular integrals allows for several simplifications to this general approach. Most importantly, by studying the behavior under dimensional shift $d\rightarrow d+2$ it is possible to reconstruct the pole and finite part of the $\varepsilon$-expansion solely in terms of known two-denominator integrals. Simultaneously, the dimensional shift can be used for analytic continuation to obtain finite integral representations – fit for numerical checks – for each order in the expansion. Like IBP relations and differential equations, also dimensional shift relations for angular integrals have their counterpart for loop integrals Tarasov:1996 ; Lee:2012 .
To facilitate keeping an overview of what follows, figure 1 provides a graphical overview of the steps taken to extract the $\varepsilon$-expansion of the general three-denominator angular integral with $m$ masses $I_{j_{1},j_{2},j_{3}}^{(m)}$.
@@
Figure 1: This flowchart provides an overview of the calculation of the $\varepsilon$-expansion of the general three-denominator angular integral $I_{j_{1}j_{2}j_{3}}^{(m)}$. In a first step, recursion relations (see app. C) derived from IBP relations (see sec. 2) are used for a reduction to the master integral $I_{1,1,1}^{(m)}$. In a second step, the double- and triple-massive integrals are expressed in terms of massless and single-massive ones through mass reduction formulae derived from the two-point splitting lemma (see sec. 3). In a third step, a combination of a dimensional-shift identity, relating integrals in $d$ and $d+2$ dimensions, with the recursion relations allows for the determination of pole and finite part in terms of known two-denominator integrals (see sec. 4). In a final step, the order $\varepsilon$ contribution is calculated by applying the method of differential equations – requiring suitable differential operators for angular integrals (see sec. 5) and again making use of the recursion relations – to the massless and single-massive master integral in $d=6-2\varepsilon$ dimensions (see sec. 6 and 7). Results are found in sec. 8. Image created with JaxoDraw Binosi:2003 .
At this point we like to briefly comment that the most common way for the analytic integration of phase space integrals is the method of reversed unitarity Anastasiou:2002 ; Anastasiou:2003 ; Anastasiou:2004 . This directly translates phase-space integrals to loop integrals, opening up the toolkit developed for these including IBP reduction and the derivation of differential equations, e.g. FIRE Smirnov:2019 , KIRA Klappert:2020 , and LiteRed Lee:2013 . In our calculation we will however perform the procedure of IBP reduction and derivation of differential equations directly on the level of angular integrals without a mapping to loop integrals, showcasing that the introduction of auxiliary cut propagators is not needed for the specific phase-space integrals under consideration. Since the integral family has a sufficiently simple structure it is feasible to proceed without the aforementioned powerful machinery. For an example of the treatment of very similar integrals with reversed unitarity technology we refer to Liu:2024 , for an example of a class of phase-space integrals which inherently require extensions of reversed unitarity to perform the IBP reduction see Baranowski:2022 .
@@
The remainder of the paper is organized as follows. In section 2, we derive IBP relations to reduce $I_{j_{1},j_{2},j_{3}}$ to master integrals. Section 3 recalls the two-point splitting lemma, which allows us to express double- and triple-massive integrals in terms of the single-massive integral. The dimensional-shift identity introduced in section 4 allows us to express the pole and finite part of the three-denominator integral solely in terms of known two-denominator integrals. In section 5, we determine differential operators, which we subsequently use to derive differential equations for the master integrals in section 6. The integration of these differential equations is discussed in section 7. Results for the massless, single-massive, double-massive, and triple-massive master integrals $I^{(0,1,2,3)}_{1,1,1}$ are given in section 8. The generalization towards angular integrals with more than three denominators is briefly discussed in section 9, before we conclude in section 10. Appendix A lists the scalar products used for the mass reduction formula, while appendices B and C give the full form of the IBP relations and the recursion relations obtained from the latter. In appendix D, a proof of the general dimensional-shift identity for angular integrals is provided. Appendix E lists the expansions for the dimensionally-shifted two-denominator integrals that are used as input for the calculation of the three-denominator master integrals. The explicit calculation of the boundary value for the massless master integral is given in appendix F. Integrals useful for the calculation of the master integrals are found in appendix G, and appendix H recalls the Clausen function appearing in the order $\varepsilon$ of the three-denominator angular integrals. The ancillary Mathematica file provides the function AngularIntegral[{j1_,j2_,j3_},{v12_,v13_,v23_,v11_,v22_,v33_}] which evaluates the three-denominator angular integral in $d=4-2\varepsilon$ dimensions up to order $\varepsilon$.
2 IBP relations
@@
Since they are the foundation for everything that follows, we start by deriving IBP relations to reduce $I_{j_{1},j_{2},j_{3}}$ to master integrals. IBP relations for angular integrals are derived in a similar fashion to loop integrals. The only difference is that permissible vectors in the IBP relation have to be tangential to the spherical integration surface. Concretely, the IBPs read
@@
where the spatial vector $\vec{\xi}$ must satisfy $\vec{\xi}\cdot\vec{k}\overset{!}{=}0$.
@@
For general $j_{1,2,3}$ these IBP relations can be combined into identities that only either raise or lower the sum of indices. This allows to bring them in the form of recursion relations that systematically reduce all integrals $I_{j_{1},j_{2},j_{3}}$ into the master integrals $I_{1,1,1}$, $I_{1,1,0}$, $I_{1,0,0}$, $I_{0,0,0}$ (+ permutation of indices), bypassing the need for Laporta’s algorithm Laporta:2000 . The explicit recursion relations are given in appendix C.
To calculate the master integrals, we use the method of differential equations. For this, we write the master integrals in the form of eq. (1) and differentiate both sides with respect to the kinematic invariants $v_{ij}$, which will lead to new angular integrals on the right-hand side – details of the differentiation are discussed in section 5. By reducing the right-hand side to master integrals via the recursion relations, we obtain a system of differential equations for the master integrals. For a pedagogical introduction to the differential equation approach in the context of loop integrals see Henn:2014 ; Badger:2023 .
Since the $\varepsilon$-expansion of two-denominator angular integrals is known Beenakker:1988 ; Somogyi:2011 ; Lyubovitskij:2021 , the only unknown master integral in our case is $I_{1,1,1}$. By explicitly inserting the known integrals into the coupled system of differential equations, it is reduced to one closed differential equation for $I_{1,1,1}$.
@@
    By two-mass splitting we can always reduce the triple-massive integral to single-massive integrals. We will be discussing this next in section 3. This simplifies the differential equation by reducing the number of variables. The number of master integrals is further reduced since the massless one-denominator integral $I^{(0)}_{1,0,0}$ is reducible to $I^{(0)}_{0,0,0}$.
@@
    By using a dimensional-shift identity $d\rightarrow d+2$ we obtain the pole and – somewhat surprisingly – also the finite part of the three-denominator integral “for free”, i.e. in terms of known integrals with fewer denominators. We will look at this in section 4.
@@
    By symmetry it suffices to look at a single derivative, say ${\partial}/{\partial v_{12}}$, for the massless case. From there we only need a single boundary value for which we choose the symmetric point $v_{12}=v_{13}=v_{23}=1$. To construct the single-massive integral from the differential equation with respect to ${\partial}/{\partial v_{11}}$ we can use the previously calculated massless integral as a boundary condition. This will be done in sections 6 and 7.
3 Mass reduction via two-point splitting
It is a truth universally acknowledged, that a single-massive angular integral is sufficient to calculate multi-mass angular integrals. This is possible due to the two-point splitting lemma Lyubovitskij:2021 ; Wunder:2024 , which states that for any two vectors $v_{1}$ and $v_{2}$, we can choose any scalar $\lambda_{(12)}$ and construct the linear combination $v_{(12)}\equiv(1-\lambda_{(12)})\,v_{1}+\lambda_{(12)}\,v_{2}$} to obtain
@@ 3 Mass reduction via two-point splitting
If we consider two massive vectors $v_{1}$ and $v_{2}$ and choose $\lambda_{(12)}$ such that $v_{(12)}$ is massless, i.e. $v_{(12)(12)}=0$, eq.(10) splits the product of two massive denominators into the sum of two single-massive denominator products. Iterative use of this identity on pairs of massive denominators allows for a reduction to the single-massive case. Hence angular integrals with an arbitrary number of masses can always be expressed through at most single-massive integrals.  [^1] A quite similar idea of splitting loop integrals into integrals with fewer masses was already put forward a long time ago by ’t Hooft and Veltman tHooft:1978 .
@@
with $X_{12}\equiv v_{12}^{2}-v_{11}v_{22}$. Note that there is the geometrical interpretation $X_{12}=4\,\mathbb{V}(\vec{v}_{1},\vec{v}_{2})^{2}+(\vec{v}_{1}-\vec{v}_{2})^{2}$, where $\mathbb{V}(\vec{v}_{1},\vec{v}_{2})$ denotes the area of the triangle spanned by $\vec{v}_{1}$ and $\vec{v}_{2}$. Hence, for real vectors $\vec{v}_{1,2}$, $X_{12}$ is non-negative and vanishes if and only if $\vec{v}_{1}=\vec{v}_{2}$. The scalar products of the auxiliary vectors are $v_{(12)(12)}=v_{(21)(21)}=0$}, $v_{1(12)}=-\lambda_{(12)}\sqrt{X_{12}}$}, $v_{2(21)}=-\lambda_{(21)}\sqrt{X_{12}}$}, and $v_{(12)(21)}=2X_{12}/(2v_{12}-v_{11}-v_{22})$}. Note that $v_{1(12)}$ respectively $v_{2(21)}$ vanish in the massless limits $v_{11}\rightarrow 0$ respectively $v_{22}\rightarrow 0$ . Using these scalar products, a compact form of the two-mass splitting is
@@
$I_{1,1,1}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})=\frac{1}{\sqrt{X_{12}}}\left[v_{(12)(21)}I_{1,1,1}^{(0)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)}\right)\right$.
$\qquad\qquad\left.-v_{1(12)}I_{1,1,1}^{(1)}\left(v_{13},v_{1(12)},v_{3(12)},v_{11}\right)-v_{2(21)}I_{1,1,1}^{(1)}\left(v_{23},v_{2(21)},v_{3(21)},v_{22}\right)\right]$ (15)
@@
$I_{1,1,1}^{(3)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33})=\frac{1}{\sqrt{X_{12}}}\Big\{v_{(12)(21)}I_{1,1,1}^{(1)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)},v_{33}\right)$
$-\frac{v_{1(12)}}{\sqrt{X_{13}}}\Big[v_{(13)(31)}I_{1,1,1}^{(0)}\left(v_{(12)(13)},v_{(12)(31)},v_{(13)(31)}\right)$
$\qquad-v_{1(13)}I_{1,1,1}^{(1)}\left(v_{1(12)},v_{1(13)},v_{(12)(13)},v_{11}\right)-v_{3(31)}I_{1,1,1}^{(1)}\left(v_{3(12)},v_{3(31)},v_{(12)(31)},v_{33}\right)\Big]$
$-\frac{v_{2(21)}}{\sqrt{X_{23}}}\Big[v_{(23)(32)}I_{1,1,1}^{(0)}\left(v_{(21)(23)},v_{(21)(32)},v_{(23)(32)}\right)$
$\qquad-v_{2(23)}I_{1,1,1}^{(1)}\left(v_{2(21)},v_{2(23)},v_{(21)(23)},v_{22}\right)-v_{3(32)}I_{1,1,1}^{(1)}\left(v_{3(21)},v_{3(32)},v_{(21)(32)},v_{33}\right)\Big]\Big\}\,$, (16)
@@
$\tilde{I}^{(m)}_{1,1,1}(v_{ij},v_{ik},v_{jk},v_{ii},v_{jj},v_{kk})\equiv\sqrt{Y_{ijk}}\,I^{(m)}_{1,1,1}(v_{ij},v_{ik},v_{jk},v_{ii},v_{jj},v_{kk})\,$, (21)
@@
$\tilde{I}_{1,1,1}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})=$ $\tilde{I}_{1,1,1}^{(0)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)}\right)-\tilde{I}_{1,1,1}^{(1)}\left(v_{13},v_{1(12)},v_{3(12)},v_{11}\right)$
$-\tilde{I}_{1,1,1}^{(1)}\left(v_{23},v_{2(21)},v_{3(21)},v_{22}\right),$ (22)